• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Ole Hansen

    Log entry time 23:41:37 on March 06, 2012

    Entry number 364262

    keyword=Hall Status

    Yves called around 23:00, reporting that the beam makes it just fine to the BPM after the chicane and in front of the target, but efforts have failed so far to see the beam on the dump. Instead, ion chambers near the target trip the beam around 300 nA. We are calling in target experts to investigate possible issues. For example, we are not 100% sure whether the target polarity is correct.